At noon, a passenger ship and a freighter left a port traveling in opposite directions. By midnight, the passenger ship was 3 times farther from port than was the freighter. How far (in mi) was the freighter from port if the distance between the ships was 88 miles?
Answer by ikleyn(52775)

.

x + 3x = 88


4x     = 88


 x     = 88/4 = 22.


ANSWER.  22 miles.